OpenStudy (anonymous):

Solve 100 ÷ 5 × 4 + 4 squared 3 ( 4 and a little 3 on top ). 69 144 0.3 1.2

OpenStudy (jdoe0001):

\(\large 100\div 5\times 4+4^23(4^3)\ ? \)

OpenStudy (jdoe0001):

either way, recall your \(\huge \mathbb{PEMDAS}\)

OpenStudy (anonymous):

?? @jdoe0001

OpenStudy (jdoe0001):

so... is \(\large 100\div 5\times 4+4^23(4^3)\) what you have?

OpenStudy (anonymous):

that's all @jdoe0001